OpenStudy (anonymous):

five candidates- A,b,C,D and E were in an election. Candidate A was elected with 30 votes, B placed second and E was in last with 3 votes. If 59 votes were cast and no two candidates had the same number of votes, what is the smallest number of votes B could have recieved?

OpenStudy (anonymous):

A = 30 B = ? C = ? D = ? E = 3 A + B + C + D + E = 59 B + C + D = 26 26 / 3 = 8.6 repeated 26 = 10 + 9 + 7 B = 10

OpenStudy (anonymous):

thanx alot